OpenStudy (anonymous):

solve ln(x-2)-ln(x+1)=lnx

OpenStudy (anonymous):

Would you like help on this?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Alright cool! So do you have an idea of what we should do first?

OpenStudy (anonymous):

would you have to equal the (x-2) and (x+1) to each other ?

OpenStudy (anonymous):

unfortunately they are logs so we can't really do so. But we can take advantage of the fact that they are in the ln's. Do you know how to get rid of the lns?

OpenStudy (anonymous):

no :c im actually extremely lost

OpenStudy (anonymous):

haha so you need to put both sides as the exponents of e, does that make sense?

OpenStudy (anonymous):

like so: \[\ln(x-2)-\ln(x+1) = lnx\]\[e^{\ln(x-2)-\ln(x+1)} =e^{ lnx}\] does this seem familiar?

OpenStudy (anonymous):

no :C ive actually never done a question like this one. my teacher posted this up as extra credit and i would appreciate it sooo much if you guided me step by step with explanation on how to do it ..

OpenStudy (anonymous):

ahhh nice extra credit! That's always a good thing. Sure, but to do this problem you will need to learn some things, is that ok with you?

OpenStudy (anonymous):

yeah i have no problem, i am in ap calculus so soon he will bring this question into topic. i am honestly not very good in calc. i end up forgetting everything the next day :( and thats my issue. but yes, i am so for it teach me everything

OpenStudy (anonymous):

oooh calc, that was a long time ago but it was enjoyable. Are you doing AB or BC?

OpenStudy (anonymous):

AB

OpenStudy (anonymous):

Anyways sorry for getting off topic...ok so you need to know this identity: \[10^{\log(x)} = x = log(10^{x})\] does this make sense?

OpenStudy (anonymous):

kind of .. yeah

OpenStudy (anonymous):

alright so just keep that in your mind btw ln is just that but instead of 10 it's for e, does that make sense?

OpenStudy (anonymous):

oh okay so ln replaces the 10 with e ? and e and 10 are the same ??

OpenStudy (anonymous):

no e is called the natural base, it's a really important number in the world. ln is the natural logarithm.

OpenStudy (anonymous):

if you have a scientific or graphing calculator it should have that value on it

OpenStudy (anonymous):

ooh okay

OpenStudy (anonymous):

ok so to show you how that works: \[e^{lnx} = x\] make sense?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so \[e^{ln(x-2)-ln(x+1)}= e^{ln(x)}\] simplifies down to so \[e^{ln(x-2)-ln(x+1)}= x\] does that makes sense?

OpenStudy (anonymous):

oooh okay yes got it, continue.

OpenStudy (anonymous):

alrighty now we need to know about another rule, so here it is: \[x^{a-b} = \frac{x^{a}}{x^{b}}\] have you seen this before?

OpenStudy (anonymous):

no i have not.

OpenStudy (anonymous):

Does it make sense though?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

alright cool so then our:\[e^{\ln(x-2)-\ln(x+1)} = x\] becomes this:\[\frac{e^{\ln(x-2)}}{e^{\ln(x+1)}} = x\] make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but then it simplifies down further because we got the whole e^ln thing: \[\frac{x-2}{x+1} = x\] do you know how to solve it now?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

well we then move the (x+1) to the other side so it becomes: \[x-2 = x(x+1)\] which is \[x-2 = x^{2} + x\] and it becomes \[-2 = x^{2}\] which we can solve to be \[x= \pm i \sqrt{2}\] does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool and that's the problem! So not that bad right?

OpenStudy (anonymous):

no its not :o easier than i thought it would be. once you presented me with the formulas i was able to understand how they connected. i also need help in one more question, if you dont mind.

OpenStudy (anonymous):

uhh sure, I might be a little slow in responding tho if that's alright with you?

OpenStudy (anonymous):

yeah its cool :) \[(1000)^x+2(100)^x-3(10)=0\]

OpenStudy (anonymous):

I am assuming you have no idea how to start this?

OpenStudy (anonymous):

nope .. :x

OpenStudy (anonymous):

haha ok so do you notice something peculiar with how they gave you the question?

OpenStudy (anonymous):

it decreases ? going fromm 1000 to 100 to 10 and equaling to 0 ?

OpenStudy (anonymous):

haha well there's that but aren't 1000, 100, and 10 all powers of 10?

OpenStudy (anonymous):

haha im just laughing at myself

OpenStudy (anonymous):

yes you are right haha

OpenStudy (anonymous):

so why don't we put everything as a power of 10, do you know how to do so?

OpenStudy (anonymous):

ummm replacing the x with 10 ??

OpenStudy (anonymous):

if i replace x with 10 the answer would be a really long number

OpenStudy (anonymous):

no all you do is this: \[(1000)^{x}+2(100)^{x}-3(10)=0\] \[10^{3^{x}}+ 2(10^{2^{x}})-3(10) = 0\] make sense?

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

now all we do is the same thing we did to the previous problem but instead of making everything the power of e, we simply log both sides and use the identity I showed you previously. Does this make sense?

OpenStudy (anonymous):

so you log both sides ? and which identity ?

OpenStudy (anonymous):

the whole: \[x^{a-b}=\frac{x^{a}}{x^{b}}\]

OpenStudy (anonymous):

part do you see it?

OpenStudy (anonymous):

oooh okay so i log both sides and then plug it into that ? or ..

OpenStudy (anonymous):

well that's how you separate the whole mess you'll have when you log both sides...oh I see, I'm sorry, that identity is the wrong one...uhhh let see...

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

uhh let me think about it for a little bit, I am like helping multiple people so I get confused every once in awhile...

OpenStudy (anonymous):

its okay haha but was that identity incorrect in general ?

OpenStudy (anonymous):

no it's a correct identity you just would use it in this case lol.

OpenStudy (anonymous):

not

OpenStudy (anonymous):

is what I mean...

OpenStudy (anonymous):

ok I got...sorry about that!

OpenStudy (anonymous):

let's start over ok?

OpenStudy (anonymous):

its okay :)

OpenStudy (anonymous):

alright so we start off with this right? (1000)^x+2(100)^x - 3(10) = 0 what we want to do is actually is divide both sides by 10: (100)^x + 2(10)^x-3 = 0 does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright now we can just say that z = 10^x, do you understand we can just arbitrarily do this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok cool so if you use that we can substitute our previous equation thusly: x^2 + 2x -3 = 0, do you get how I got this? from this: 100^x+2(10^x)-3 = 0?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

cool and now we can just solve that quadratic right? so what is the solution to that quadratic?

OpenStudy (anonymous):

(x+3)(x-1)=0 ?

OpenStudy (anonymous):

yup so x = -3,1. But wait!!! That's not x right? that's actually z! so z = -3,1 but we want x, so we go back to how we defined z: z = 10^x or x = log(z)...Therefore x = log(-3),log(1). But wait! we can't have a log of a negative number of log(-3) doesn't make any sense. And log(1) is equal to 10. So now we have our final solution of x = 10! Does this all make sense?

OpenStudy (anonymous):

where did z come from though ?

OpenStudy (anonymous):

remember we arbitrarily defined previous that z = 10^x? Like a few comments ago.

OpenStudy (anonymous):

oooh yes yes sorry

OpenStudy (anonymous):

so does everything make sense?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

nice! sorry for leading you astray! But I'm glad that you understood everything!

OpenStudy (anonymous):

thank you so much !! have a good night :)

OpenStudy (anonymous):

you as well!